OpenStudy (vishweshshrimali5):

Let N be the integer next above \((\sqrt{3} + 1)^{2014}\). The greatest integer p such that 16^p divides N is _____.

OpenStudy (vishweshshrimali5):

Okay I am pretty sure that I am struck again. I cannot figure it out that why does this always happen with me ?? Its really frustrating...

OpenStudy (vishweshshrimali5):

I am going to do something about this N first

OpenStudy (vishweshshrimali5):

First of all for finding N, I can write N as: N = [\(\sqrt{3} + 1)^{2014}\)] + 1 Where, [x] means greatest integer function

OpenStudy (vishweshshrimali5):

Now I have to find out [\(\sqrt{3} + 1)^{2014}\)]. Okay, so, uhm, [(\(\sqrt{3} + 1)^{2014}\)] = (\(\sqrt{3} + 1)^{2014}\) - {(\(\sqrt{3} + 1)^{2014}\)} Where, {x} represents fractional part of x.

OpenStudy (vishweshshrimali5):

Now, I know this, \(\sqrt{3} = 1.732\) So, I can write it as 1 + some pure fraction (let f). So, I have, \(\sqrt{3} = 1+f\) So, \((\sqrt{3} + 1)^{2014} = (f+2)^{2014}\)

OpenStudy (vishweshshrimali5):

Okay so I stuck again. And I am going to wait with my hands folded and cursing maths in my mind until something else appears in my mind.

OpenStudy (vishweshshrimali5):

Are you guys doing the same too ? lol

OpenStudy (anonymous):

yes :))

OpenStudy (vishweshshrimali5):

Okay, lets do it together.

OpenStudy (vishweshshrimali5):

Okay I think there is something better than that: \(\large{(\sqrt{3} + 1)^{2014} = {(\sqrt{3})^{2014}}_i + {(\sqrt{3})^{2013}}_f + {(\sqrt{3})^{2012}}_i + ... + {(\sqrt{3})^{0}}_i}\)

OpenStudy (vishweshshrimali5):

I have put \(i\) to represent that the term is integer and \(f\) to represent that the term is fractional.

OpenStudy (vishweshshrimali5):

Now, I know for sure, that integer will come from both the terms. So, I am going to simplify only the odd terms.

OpenStudy (vishweshshrimali5):

*odd powered terms

OpenStudy (vishweshshrimali5):

Now, I know that, \(\sqrt{3} = 1 + f\) where, f<1 So, for some odd power of (\(\sqrt{3})\), say, 2n - 1, I am going to write its expansion.

OpenStudy (vishweshshrimali5):

\(\large{(\sqrt{3})^{2n-1} = (1 + f)^{2n-1}}\)

OpenStudy (vishweshshrimali5):

I think its safe to assume that: \(N = (\sqrt{3} + 1)^{2014} + (\sqrt{3} - 1)^{2014}\) As, I know that \((\sqrt{3} + 1)^{2014}\) is going to sum of terms with "i" which would be integer and sum of terms with "f" which would be fractional. Now, for N is an integer, I must add some positive number to it to make it an integer and that number to be added must be the smallest possible number which in my view is \((\sqrt{3} - 1)^{2014}\), as it removes all the fractional terms (i.e. terms marked as "f").

OpenStudy (vishweshshrimali5):

Now, \[\large{(\sqrt{3} + 1)^{2014} + (\sqrt{3} - 1)^{2014}}\] \[\large{= 2^{1007}[(2+\sqrt{3})^{1007} + (2-\sqrt{3})^{1007}]}\] \[\large{= 2^{1009}[2^{1006} + \left(\begin{matrix}1007 \\ 2\end{matrix}\right)*2^{1004}*3 + ... + \left(\begin{matrix}1007 \\ 1006\end{matrix}\right)*3^{503}]}\] Now, the part in [ ] brackets is not divisible by . So, \[\large{2^{1009} = 2*2^{1008} = 2* 16^{252}}\] So, required value of p = 252.

OpenStudy (vishweshshrimali5):

*sigh* Its done !!!!!!!!

mathslover (mathslover):

o.O you wrote LaTeX ?

OpenStudy (vishweshshrimali5):

Hey I resent that statement ?

OpenStudy (vishweshshrimali5):

Well, uhm, sort of.

OpenStudy (vishweshshrimali5):

Now, the only thing remains is getting this solution checked. Suggestions for who to ask for help ?

mathslover (mathslover):

@nipunmalhotra93 @dan815 may help you!

OpenStudy (vishweshshrimali5):

I am going to add some tags from my side as well. @BSwan , @ganeshie8 , @mukushla , @mathmale .

OpenStudy (vishweshshrimali5):

I just want to know if my answer is correct or not. I only have question and my approach to the problem. I don't have the answer.

OpenStudy (vishweshshrimali5):

Well I have to go to have dinner. Sorry friends. Bye

OpenStudy (anonymous):

quite right :-) you nailed it

OpenStudy (vishweshshrimali5):

Thanks @mukushla

OpenStudy (anonymous):

no problem

ganeshie8 (ganeshie8):

wow ! the mighty mukushla :)

OpenStudy (vishweshshrimali5):

Hail mukushla lol

OpenStudy (anonymous):

lol, gane vish solved it not me :-)

OpenStudy (vishweshshrimali5):

;)

OpenStudy (vishweshshrimali5):

Hey its not "pick on vishu" day.. :P

OpenStudy (anonymous):

:p

OpenStudy (vishweshshrimali5):

good night guys. Its time to go to bed. :)

OpenStudy (vishweshshrimali5):

Enjoy the life 'coz its worth enjoying !!